# Global well-posedness of the elastic-viscous-plastic sea-ice model with the inviscid Voigt-regularisation

### v1

@@ 2 Preliminaries
$A\lesssim B.$ (2.1)
We recall a few results that will be used in this paper. We first recall the Brézis-Gallouët–Wainger inequality, which was originally proved in [13, 12]:
@@
Let $v\in H^{2}(\mathbb{T}^{2})$ (where $v$ is dimensionless), then it satisfies the following inequality
$\lVert v\rVert_{L^{\infty}}\leq C\left(\lVert v\rVert_{H^{1}}\sqrt{\log\bigg{(}1+C_{L}\lVert v\rVert_{H^{2}}\bigg{)}}+1\right),$ (2.2)
@@
We will also use the Ladyzhenskaya inequality in two space dimensions
$\lVert f\rVert_{L^{4}}\lesssim\lVert f\rVert_{L^{2}}^{1/2}\lVert\nabla f\rVert_{L^{2}}^{1/2}+\lVert f\rVert_{L^{2}};$ (2.3)
as well as the logarithmic Grönwall inequality, which in this form can be found in [16,  Lemma 2.5]:
@@
Let $v_{1},v_{2}\in H^{2}(\mathbb{T}^{2})$ and recall that the notation $D(v)$ was defined in equation (1.1), then we have
$\left\lVert\sqrt{\lvert D(v_{1})\rvert^{2}+\epsilon^{2}}-\sqrt{\lvert D(v_{2})\rvert^{2}+\epsilon^{2}}\right\rVert_{L^{2}}\leq\left\lVert\nabla v_{1}-\nabla v_{2}\right\rVert_{L^{2}},$ (2.7)
@@
This implies inequality (2.7) for smooth $v_{1}$ and $v_{2}$. For $v_{1},v_{2}\in H^{2}(\mathbb{T}^{2})$, one can proceed by choosing smooth approximating sequences of $H^{2}(\mathbb{T}^{2})$ functions and conclude the proof. ∎
Although it is not necessary for the construction, we show in the following lemma that $\sigma$ remains symmetric under the evolution of the Voigt-EVP model if it is initially symmetric:
@@
Consider a solution $(u,\sigma)$ to either system (1.12) or (1.13), which has the regularity described in Theorem 1.2, with $\sigma_{0}\in H^{3}(\mathbb{T}^{2};\mathbb{R}^{2\times 2})$ being symmetric. That is, $u\in C([0,T];H^{2}(\mathbb{T}^{2}))$ and $\sigma\in C([0,T];H^{3}(\mathbb{T}^{2}))$ for arbitrary $T>0$. Then $\sigma(x,t)=\sigma(x,t)^{\top}$ for all $0\leq t\leq T$ and all $x\in\mathbb{T}^{2}$.
@@
$W(\sigma)\coloneqq\frac{1}{2}(\sigma-\sigma^{\top}).$ (2.8)
Then, from equation (1.12b) or equation (1.13b), one has
$\frac{1}{\mathcal{E}}\partial_{t}(I-\alpha^{2}\Delta)W(\sigma)+\frac{4H}{P}W(\sigma)=0,$ (2.9)
@@
$\sup_{t\in[0,T]}\big{[}\lVert W(\sigma)(t)\rVert_{L^{2}}^{2}+\alpha^{2}\lVert\nabla W(\sigma)(t)\rVert_{L^{2}}^{2}\big{]}\leq\lVert W(\sigma_{0})\rVert_{L^{2}}^{2}+\alpha^{2}\lVert\nabla W(\sigma_{0})\rVert_{L^{2}}^{2}=0.$ (2.10)
@@
3 Global solutions to the intermediate system (1.13)
Our strategy of establishing the global solvability of system (1.13) is the following: First, we will write down the Galerkin approximation scheme. At each level of the scheme, by directly applying the Picard theory we obtain a local-in-time solution of the Galerkin system. Second, we establish better regularity for the approximating solutions (as well as their global existence in time), and pass to the limit in the Galerkin scheme. Finally, we establish the uniqueness of the strong solution.
@@ 3.1 The Galerkin approximation scheme
$\partial_{t}u_{N}=\nabla\cdot\sigma_{N}+P_{N}\mathcal{T}_{a}+P_{N}\mathcal{T}_{w}^{N}+\Omega u^{\perp}_{N}-gP_{N}\nabla H_{0},$ (3.1a)
$\begin{gathered}\frac{1}{\mathcal{E}}\partial_{t}\big{(}\sigma_{N}-\alpha^{2}\Delta\sigma_{N}\big{)}+P_{N}\bigg{[}\frac{4\mathcal{D}_{\epsilon}^{N}}{P}(\sigma_{N}-\frac{1}{2}\operatorname{Tr}\sigma_{N}\mathbb{I}_{2})\bigg{]}+P_{N}\bigg{[}\frac{\mathcal{D}_{\epsilon}^{N}}{2P}\operatorname{Tr}\sigma_{N}\mathbb{I}_{2}\bigg{]}\\ +\frac{P_{N}\mathcal{D}_{\epsilon}^{N}}{2}\mathbb{I}_{2}=D(u_{N}),\end{gathered}$ (3.1d)
$u\lvert_{t=0}=P_{N}u_{0},\quad\sigma\lvert_{t=0}=P_{N}\sigma_{0},$ (3.1e)
where $P_{N}$ is the projection onto the Fourier modes up to order $N$, including the zeroth mode. In system (3.1), we have used the notation
$\mathcal{T}_{w}^{N}$ $=c_{w}\rho_{w}\lvert U_{w}-u_{N}\rvert\bigg{[}(U_{w}-u_{N})\cos\theta+(U_{w}-u_{N})^{\perp}\sin\theta\bigg{]}$ (3.2)
@@
$\mathcal{D}_{\epsilon}^{N}$ $=\sqrt{\lvert D(u_{N})\rvert^{2}+\epsilon^{2}}.$ (3.3)
Notice that system (3.1) is a first-order system of ordinary differential equations for $\{\hat{u}_{k},\hat{\sigma}_{k}\}_{|k|\leq N}$, with a locally Lipschitz continuous vector field with respect to the unknowns. Thus, the Picard theory implies that, for each initial datum $(P_{N}u_{0},P_{N}\sigma_{0})$, there exists a unique local-in-time solution.
@@ 3.2 Uniform-in-$(N,\epsilon)$ estimates
To establish the uniform-in-$N$ estimates for the solutions of system (3.1), let
$\tau_{N}\coloneqq\sigma_{N}+\frac{P}{2}\mathbb{I}_{2}.$ (3.4)
Recall that the pressure $P>0$ is constant. We then rewrite equation (3.1d) as follows:
$\begin{gathered}\frac{1}{\mathcal{E}}\partial_{t}\bigg{(}\tau_{N}-\alpha^{2}\Delta\tau_{N}\bigg{)}+P_{N}\bigg{[}\frac{4\mathcal{D}_{\epsilon}^{N}}{P}\bigg{(}\tau_{N}-\frac{1}{2}\operatorname{Tr}\tau_{N}\mathbb{I}_{2}\bigg{)}\bigg{]}\\ +P_{N}\bigg{[}\frac{\mathcal{D}_{\epsilon}^{N}}{2P}\operatorname{Tr}\tau_{N}\mathbb{I}_{2}\bigg{]}=D(u_{N}).\end{gathered}$ (3.5)
$L^{2}$-estimate. We start by computing the $L^{2}(\mathbb{T}^{2})$ estimates for the solution of system (3.1) on its maximal (positive) time interval of existence. Taking the $L^{2}$-inner product of (3.1a) and (3.5) with $u_{N}$ and $\tau_{N}$, respectively, applying integration by parts, and summing the results yields
$\begin{gathered}\frac{1}{2}\frac{d}{dt}\bigg{[}\lVert u_{N}\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\|\tau_{N}\|_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\|\nabla\tau_{N}\|_{L^{2}}^{2}\biggr{]}\\ +\int_{\mathbb{T}^{2}}\biggl{[}\frac{4\mathcal{D}_{\epsilon}^{N}}{P}|\tau_{N}-\frac{1}{2}\operatorname{Tr}\tau_{N}\mathbb{I}_{2}|^{2}+\frac{\mathcal{D}_{\epsilon}^{N}}{2P}|\operatorname{Tr}\tau_{N}|^{2}\biggr{]}\,\ {\rm d}{x}\\ =\int_{\mathbb{T}^{2}}\bigg{[}\mathcal{T}_{a}+\mathcal{T}_{w}^{N}+\Omega u^{\perp}_{N}-g\nabla H_{0}\bigg{]}\cdot u_{N}\ {\rm d}{x}\leq C+C\|u_{N}\|_{L^{2}}^{2},\end{gathered}$ (3.6)
@@
where we use the fact that $\cos\theta\geq 0$ since $\theta\in[0,\frac{\pi}{4}]$. Notice that the constant $C\in(0,\infty)$ in (3.6) depends only on $\Omega,g,P,H_{0},U_{a},U_{w},c_{a},c_{w},\phi,\theta,\rho_{a},$ and $\rho_{w}$, but not on $\epsilon$ and $N$. Moreover, we also have used the following cancellation, which is similar in structure to the Oldroyd-B model,
$\int_{\mathbb{T}^{2}}\bigg{[}u_{N}\cdot(\nabla\cdot\sigma_{N})+\tau_{N}:D(u_{N})\bigg{]}\ {\rm d}{x}=0.$ (3.7)
Therefore, applying the Grönwall inequality to estimate (3.6) on a time interval of existence $[0,T]$ of the approximating Galerkin solution leads to
$\sup_{t\in[0,T]}\bigg{[}\lVert u_{N}(t)\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\left\lVert\sigma_{N}(t)+\frac{P}{2}\mathbb{I}_{2}\right\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\left\lVert\nabla\bigg{[}\sigma_{N}(t)+\frac{P}{2}\mathbb{I}_{2}\bigg{]}\right\rVert_{L^{2}}^{2}\bigg{]}\leq Ce^{CT},$ (3.8)
uniformly in $N$. Since the right-hand side remains finite for any $T<\infty$, we conclude that the time interval of existence of the approximating Galerkin solutions is $[0,\infty)$ and estimate (3.8) therefore holds for any $T<\infty$.
@@
$V_{w}:=U_{w}-u_{N}.$ (3.9)
Recalling (3.2), one can write
$\begin{gathered}-\int_{\mathbb{T}^{2}}P_{N}\mathcal{T}_{w}^{N}\cdot\Delta u_{N}\ {\rm d}{x}=\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}P_{N}\partial_{i}\mathcal{T}_{w}^{N}\cdot\partial_{i}u_{N}\ {\rm d}{x}=\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\partial_{i}\mathcal{T}_{w}^{N}\cdot\partial_{i}u_{N}\ {\rm d}{x}\\ =\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\partial_{i}\mathcal{T}_{w}^{N}\cdot\partial_{i}{U_{w}}\ {\rm d}{x}-\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}c_{w}\rho_{w}\bigl{[}\partial_{i}(\lvert V_{w}\rvert V_{w})\cos\theta\bigr{]}\cdot\partial_{i}V_{w}\ {\rm d}{x}\\ -\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}c_{w}\rho_{w}\partial_{i}(|V_{w}|V_{w}^{\perp})\cdot\partial_{i}V_{w}\sin\theta\ {\rm d}{x}=:I_{1}+I_{2}+I_{3}.\end{gathered}$ (3.10)
For $i=1,2$, one can calculate
$\partial_{i}(|V_{w}|V_{w})=|V_{w}|\partial_{i}V_{w}+\dfrac{V_{w}\cdot\partial_{i}V_{w}}{|V_{w}|}V_{w}.$ (3.11)
@@
$|I_{1}|\lesssim\|V_{w}\|_{L^{2}}\|\nabla V_{w}\|_{L^{2}}\lesssim 1+\|{u_{N}}\|_{L^{2}}^{2}+\|{\nabla u_{N}}\|_{L^{2}}^{2},$ (3.12)
$|I_{3}|\leq c_{w}\rho_{w}|\sin\theta|\int_{\mathbb{T}^{2}}|V_{w}||\nabla V_{w}|^{2}\ {\rm d}{x}.$ (3.13)
Meanwhile, substituting (3.11) in $I_{2}$, one can compute
$I_{2}=-c_{w}\rho_{w}\cos\theta\int_{\mathbb{T}^{2}}\left(|V_{w}||\nabla V_{w}|^{2}+\frac{\left|\nabla\big{(}|V_{w}|^{2}\big{)}\right|^{2}}{|V_{w}|}\right)\ {\rm d}{x}.$ (3.14)
Consequently, (3.10)–(3.14) leads to
$-\int_{\mathbb{T}^{2}}P_{N}\mathcal{T}_{w}^{N}\cdot\Delta u_{N}\ {\rm d}{x}=\int_{\mathbb{T}^{2}}\nabla\mathcal{T}_{w}^{N}:\nabla u_{N}\lesssim 1+\|{u_{N}}\|_{L^{2}}^{2}+\|{\nabla u_{N}}\|_{L^{2}}^{2}$ (3.15)
@@
Taking the $L^{2}$-inner product of (3.1a) and (3.1d) with $-\Delta u_{N}$ and $-\Delta\sigma_{N}$, respectively, applying integration by parts, and summing up the results leads to the following estimate
$\frac{1}{2}\frac{d}{dt}\bigg{[}\lVert\nabla u_{N}\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\nabla\sigma_{N}\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\lVert\Delta\sigma_{N}\rVert_{L^{2}}^{2}\bigg{]}$

### v2

@@ 2 Preliminaries
$A\lesssim B.$ (15)
We recall a few results that will be used in this paper. We first recall the Brézis-Gallouët–Wainger inequality, which was originally proved in Refs. [12, 13]:
@@
Let $v\in H^{2}(\mathbb{T}^{2})$ (where $v$ is dimensionless, i.e. it does not have physical units), then it satisfies the following inequality
$\lVert v\rVert_{L^{\infty}}\leq C\left(\lVert v\rVert_{H^{1}}\sqrt{\log\bigg(1+C_{L}\lVert v\rVert_{H^{2}}\bigg)}+1\right),$ (16)
@@
We will also use the Ladyzhenskaya inequality in two space dimensions. Let $f\in H^{1}(\mathbb{T}^{2})$, then we have
$\lVert f\rVert_{L^{4}}\lesssim\lVert f\rVert_{L^{2}}^{1/2}\lVert\nabla f\rVert_{L^{2}}^{1/2}+\lVert f\rVert_{L^{2}}.$ (17)
In addition, we recall the logarithmic Grönwall inequality, which in this form can be found in Lemma 2.5 in Ref. [16]:
@@
Let $v_{1},v_{2}\in H^{2}(\mathbb{T}^{2})$ and recall the notation $D(v)$ defined in Eq. (1), then we have
$\left\lVert\sqrt{\lvert D(v_{1})\rvert^{2}+\epsilon^{2}}-\sqrt{\lvert D(v_{2})\rvert^{2}+\epsilon^{2}}\right\rVert_{L^{2}}\leq\left\lVert\nabla v_{1}-\nabla v_{2}\right\rVert_{L^{2}},$ (21)
@@
This implies inequality (21) for smooth $v_{1}$ and $v_{2}$. For $v_{1},v_{2}\in H^{2}(\mathbb{T}^{2})$, one can proceed by choosing smooth approximating sequences of $H^{2}(\mathbb{T}^{2})$ functions and conclude the proof. ∎
We show in the following lemma that $\sigma$ remains symmetric under the evolution of the Voigt-EVP model if it is initially symmetric:
@@
Consider a solution $(u,\sigma)$ to either system (12) or (13), which has the regularity described in Theorem 1.2, with $\sigma_{0}\in H^{3}(\mathbb{T}^{2};\mathbb{R}^{2\times 2})$ being symmetric. That is, $u\in C([0,T];H^{2}(\mathbb{T}^{2}))$ and $\sigma\in C([0,T];H^{3}(\mathbb{T}^{2}))$ for arbitrary $T>0$. Then $\sigma(x,t)=\sigma(x,t)^{\top}$ for all $0\leq t\leq T$ and all $x\in\mathbb{T}^{2}$.
@@
$W(\sigma)\coloneqq\frac{1}{2}(\sigma-\sigma^{\top}).$ (22)
Then, from Eq. (12b) or Eq. (13b), one has
$\frac{1}{\mathcal{E}}\partial_{t}(I-\alpha^{2}\Delta)W(\sigma)+\frac{4H}{P}W(\sigma)=0,$ (23)
@@
$\sup_{t\in[0,T]}\big[\lVert W(\sigma)(t)\rVert_{L^{2}}^{2}+\alpha^{2}\lVert\nabla W(\sigma)(t)\rVert_{L^{2}}^{2}\big]\leq\lVert W(\sigma_{0})\rVert_{L^{2}}^{2}+\alpha^{2}\lVert\nabla W(\sigma_{0})\rVert_{L^{2}}^{2}=0.$ (24)
@@
3 Global solutions to the Intermediate System (13)
Our strategy of establishing the global solvability of system (13) is the following: First, we will write down the Galerkin approximation scheme. At each level of the scheme, by directly applying the Picard theory we obtain a local-in-time solution of the Galerkin system. Second, we establish better regularity for the approximating solutions (as well as their global existence in time), and pass to the limit in the Galerkin scheme using the Aubin-Lions compactness theorem. Finally, we establish the uniqueness of the strong solution.
@@ 3.1 The Galerkin approximation scheme
$\partial_{t}u_{N}=\nabla\cdot\sigma_{N}+\mathbf{P}_{N}\mathcal{T}_{a}+\mathbf{P}_{N}\mathcal{T}_{w}^{N}+\Omega u^{\perp}_{N}-g\mathbf{P}_{N}\nabla H_{0},$ (25a)
$\begin{gathered}\frac{1}{\mathcal{E}}\partial_{t}\big(\sigma_{N}-\alpha^{2}\Delta\sigma_{N}\big)+\mathbf{Q}_{N}\bigg[\frac{4\mathcal{D}_{\epsilon}^{N}}{P}(\sigma_{N}-\frac{1}{2}\operatorname{Tr}\sigma_{N}\mathbb{I}_{2})\bigg]+\mathbf{Q}_{N}\bigg[\frac{\mathcal{D}_{\epsilon}^{N}}{2P}\operatorname{Tr}\sigma_{N}\mathbb{I}_{2}\bigg]\\ +\frac{\mathbf{Q}_{N}\mathcal{D}_{\epsilon}^{N}}{2}\mathbb{I}_{2}=D(u_{N}),\end{gathered}$ (25d)
$u\lvert_{t=0}=\mathbf{P}_{N}u_{0},\quad\sigma\lvert_{t=0}=\mathbf{Q}_{N}\sigma_{0},$ (25e)
where $\mathbf{P}_{N}$ is the projection of two-dimensional vector fields onto the Fourier modes up to order $N$, including the zeroth mode, while $\mathbf{Q}_{N}$ is the projection of $2\times 2$ (symmetric) matrices onto the Fourier modes up to order $N$. In system (25), we have used the notation
$\mathcal{T}_{w}^{N}$ $=c_{w}\rho_{w}\lvert U_{w}-u_{N}\rvert\bigg[(U_{w}-u_{N})\cos\theta+(U_{w}-u_{N})^{\perp}\sin\theta\bigg]$ (26)
@@
$\mathcal{D}_{\epsilon}^{N}$ $=\sqrt{\lvert D(u_{N})\rvert^{2}+\epsilon^{2}}.$ (27)
Notice that system (25) is a first-order system of ordinary differential equations for $\{\hat{u}_{k},\hat{\sigma}_{k}\}_{|k|\leq N}$, with a locally Lipschitz continuous vector field with respect to the unknowns. Thus, the Picard theory (see for example Ref. [28]) implies that, for each initial datum $(\mathbf{P}_{N}u_{0},\mathbf{Q}_{N}\sigma_{0})$, there exists a unique local-in-time solution.
@@ 3.2 Uniform-in-$(N,\epsilon)$ estimates
To establish the uniform-in-$N$ estimates for the solutions of system (25), let
$\tau_{N}\coloneqq\sigma_{N}+\frac{P}{2}\mathbb{I}_{2}.$ (28)
Recall that the pressure $P>0$ is constant. We then rewrite Eq. (25d) as follows:
$\begin{gathered}\frac{1}{\mathcal{E}}\partial_{t}\bigg(\tau_{N}-\alpha^{2}\Delta\tau_{N}\bigg)+\mathbf{Q}_{N}\bigg[\frac{4\mathcal{D}_{\epsilon}^{N}}{P}\bigg(\tau_{N}-\frac{1}{2}\operatorname{Tr}\tau_{N}\mathbb{I}_{2}\bigg)\bigg]\\ +\mathbf{Q}_{N}\bigg[\frac{\mathcal{D}_{\epsilon}^{N}}{2P}\operatorname{Tr}\tau_{N}\mathbb{I}_{2}\bigg]=D(u_{N}).\end{gathered}$ (29)
$L^{2}$-estimate. We start by computing the $L^{2}(\mathbb{T}^{2})$ estimates for the solution of system (25) on its maximal (positive) time interval of existence. Taking the $L^{2}$-inner product of (25a) and (29) with $u_{N}$ and $\tau_{N}$, respectively, applying integration by parts, and summing the results yields
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert u_{N}\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\|\tau_{N}\|_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\|\nabla\tau_{N}\|_{L^{2}}^{2}\biggr]\\ +\int_{\mathbb{T}^{2}}\biggl[\frac{4\mathcal{D}_{\epsilon}^{N}}{P}|\tau_{N}-\frac{1}{2}\operatorname{Tr}\tau_{N}\mathbb{I}_{2}|^{2}+\frac{\mathcal{D}_{\epsilon}^{N}}{2P}|\operatorname{Tr}\tau_{N}|^{2}\biggr]\,\ {\rm d}{x}\\ =\int_{\mathbb{T}^{2}}\bigg[\mathcal{T}_{a}+\mathcal{T}_{w}^{N}+\Omega u^{\perp}_{N}-g\nabla H_{0}\bigg]\cdot u_{N}\ {\rm d}{x}\leq C+C\|u_{N}\|_{L^{2}}^{2},\end{gathered}$ (30)
@@
where we use the fact that $\cos\theta\geq 0$ since $\theta\in[0,\frac{\pi}{4}]$. Notice that the constant $C\in(0,\infty)$ in (30) depends only on $\Omega,g,P,H_{0},U_{a},U_{w},c_{a},c_{w},\phi,\theta,\rho_{a},$ and $\rho_{w}$, but not on $\epsilon$ and $N$. We remark that estimate (30) would also hold if one considers the intermediate model with the regularised and nonsimplified strain rate $\overline{\mathcal{D}}_{\epsilon}^{N}$ instead of $\mathcal{D}_{\epsilon}^{N}$ (as both have definite sign). Moreover, we also have used the following cancellation, which is similar in structure to the Oldroyd-B model,
$\int_{\mathbb{T}^{2}}\bigg[u_{N}\cdot(\nabla\cdot\sigma_{N})+\tau_{N}:D(u_{N})\bigg]\ {\rm d}{x}=0.$ (31)
Therefore, applying the Grönwall inequality to estimate (30) on a time interval of existence $[0,T]$ of the approximating Galerkin solution leads to
$\sup_{t\in[0,T]}\bigg[\lVert u_{N}(t)\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\left\lVert\sigma_{N}(t)+\frac{P}{2}\mathbb{I}_{2}\right\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\left\lVert\nabla\bigg[\sigma_{N}(t)+\frac{P}{2}\mathbb{I}_{2}\bigg]\right\rVert_{L^{2}}^{2}\bigg]\leq Ce^{CT},$ (32)
uniformly in $N$ (and also $\epsilon$). Since the right-hand side remains finite for any $T\in(0,\infty)$, we conclude that the time interval of existence of the approximating Galerkin solutions is $0,\infty)$ and estimate (32) therefore holds for any $T\in(0,\infty)$.
@@
$V_{w}:=U_{w}-u_{N}.$ (33)
Recalling (26), one can write
$\begin{gathered}-\int_{\mathbb{T}^{2}}\mathbf{P}_{N}\mathcal{T}_{w}^{N}\cdot\Delta u_{N}\ {\rm d}{x}=\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\mathbf{P}_{N}\partial_{i}\mathcal{T}_{w}^{N}\cdot\partial_{i}u_{N}\ {\rm d}{x}=\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\partial_{i}\mathcal{T}_{w}^{N}\cdot\partial_{i}u_{N}\ {\rm d}{x}\\ =\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\partial_{i}\mathcal{T}_{w}^{N}\cdot\partial_{i}{U_{w}}\ {\rm d}{x}-\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}c_{w}\rho_{w}\bigl[\partial_{i}(\lvert V_{w}\rvert V_{w})\cos\theta\bigr]\cdot\partial_{i}V_{w}\ {\rm d}{x}\\ -\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}c_{w}\rho_{w}\partial_{i}(|V_{w}|V_{w}^{\perp})\cdot\partial_{i}V_{w}\sin\theta\ {\rm d}{x}=:I_{1}+I_{2}+I_{3},\end{gathered}$ (34)
where we used that the Galerkin projection commutes with derivatives. For $i=1,2$, one can calculate
$\partial_{i}(|V_{w}|V_{w})=|V_{w}|\partial_{i}V_{w}+\dfrac{V_{w}\cdot\partial_{i}V_{w}}{|V_{w}|}V_{w}.$ (35)
@@
$|I_{1}|\lesssim\|V_{w}\|_{L^{2}}\|\nabla V_{w}\|_{L^{2}}\lesssim 1+\|{u_{N}}\|_{L^{2}}^{2}+\|{\nabla u_{N}}\|_{L^{2}}^{2},$ (36)
$|I_{3}|\leq\left\lvert\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}c_{w}\rho_{w}\left(|V_{w}|\partial_{i}V_{w}^{\perp}+\dfrac{V_{w}\cdot\partial_{i}V_{w}}{|V_{w}|}V_{w}^{\perp}\right)\cdot\partial_{i}V_{w}\sin\theta\ {\rm d}{x}\right\rvert$
$=\left\lvert\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}c_{w}\rho_{w}\dfrac{V_{w}\cdot\partial_{i}V_{w}}{|V_{w}|}V_{w}^{\perp}\cdot\partial_{i}V_{w}\sin\theta\ {\rm d}{x}\right\rvert$ (37)
$\leq c_{w}\rho_{w}|\sin\theta|\int_{\mathbb{T}^{2}}|V_{w}||\nabla V_{w}|^{2}\ {\rm d}{x}.$ (38)
Meanwhile, substituting (35) in $I_{2}$, one can compute
$I_{2}=-c_{w}\rho_{w}\cos\theta\int_{\mathbb{T}^{2}}\left(|V_{w}||\nabla V_{w}|^{2}+\frac{1}{4}\frac{\left|\nabla\big(|V_{w}|^{2}\big)\right|^{2}}{|V_{w}|}\right)\ {\rm d}{x}.$ (39)
Consequently, (34)–(39) leads to
$\int_{\mathbb{T}^{2}}\mathbf{P}_{N}\mathcal{T}_{w}^{N}\cdot\Delta u_{N}\ {\rm d}{x}=\int_{\mathbb{T}^{2}}\nabla\mathcal{T}_{w}^{N}:\nabla u_{N}\lesssim 1+\|{u_{N}}\|_{L^{2}}^{2}+\|{\nabla u_{N}}\|_{L^{2}}^{2}$ (40)
@@
Taking the $L^{2}$-inner product of (25a) and (25d) with $\Delta u_{N}$ and $\Delta\sigma_{N}$, respectively, applying integration by parts, and summing up the results leads to the following estimate
$\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\nabla u_{N}\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\nabla\sigma_{N}\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\lVert\Delta\sigma_{N}\rVert_{L^{2}}^{2}\bigg]$
@@
$\lesssim\,$ $\lVert\mathcal{D}_{\epsilon}^{N}\rVert_{L^{2}}\lVert\sigma_{N}\rVert_{L^{\infty}}\lVert\Delta\sigma_{N}\rVert_{L^{2}}+\lVert\Delta\sigma_{N}\rVert_{L^{2}}^{2}+\lVert\nabla u_{N}\rVert_{L^{2}}^{2}+\|u_{N}\|_{L^{2}}^{2}+1$
$\lesssim\,$ $\left[\sqrt{\log(e+C_{L}R_{1}\lVert\sigma_{N}\rVert_{H^{2}})}\lVert\sigma_{N}\rVert_{H^{1}}+1\right]\big[\lVert\nabla u_{N}\rVert_{L^{2}}^{2}+\lVert\Delta\sigma_{N}\rVert_{L^{2}}^{2}\big]$
$+\lVert\Delta\sigma_{N}\rVert_{L^{2}}^{2}+\lVert\nabla u_{N}\rVert_{L^{2}}^{2}+\|u_{N}\|_{L^{2}}^{2}+1,$
$\lesssim\,$ $\left[\log(e+C_{L}R_{1}\lVert\sigma_{N}\rVert_{H^{2}})+\lVert\sigma_{N}\rVert_{H^{1}}^{2}+1\right]\big[\lVert\nabla u_{N}\rVert_{L^{2}}^{2}+\lVert\Delta\sigma_{N}\rVert_{L^{2}}^{2}\big]$
$+\lVert\Delta\sigma_{N}\rVert_{L^{2}}^{2}+\lVert\nabla u_{N}\rVert_{L^{2}}^{2}+\|u_{N}\|_{L^{2}}^{2}+1$ (42)
where we have used (40), the Brézis-Gallouët-Wainger inequality (16), as well as Young’s inequality. Note that we have introduced a constant $R_{1}>0$ in order to make the quantity $R_{1}\sigma_{N}$ dimensionless (so that we can apply the Brézis-Gallouët-Wainger inequality), and in particular $R_{1}\sigma_{N}$ does not have physical units. One can choose $R_{1}=\lVert\sigma_{0}\rVert_{L^{\infty}}^{-1}$ for example (for nonzero initial data). We observe that for the ${H}^{1}$-estimate one can also treat the nonsimplified strain rate $\overline{\mathcal{D}}_{\epsilon}^{N}$ instead of $\mathcal{D}_{\epsilon}^{N}$ by using estimate (79). In addition, we have used the following cancellation identity
$\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\bigg[\partial_{i}u_{N}\cdot(\nabla\cdot\partial_{i}\sigma_{N})+\partial_{i}\sigma_{N}:D(\partial_{i}u_{N})\bigg]\ {\rm d}x=0.$ (43)
Thanks to the $L^{2}$-estimate in (32), we can apply the logarithmic Grönwall inequality from Lemma 2.2 to (42), where we set
$A(t)\coloneqq\lVert u_{N}\rVert_{H^{1}}^{2}+\lVert\sigma_{N}\rVert_{H^{2}}^{2},\quad f(t)\coloneqq 1,\quad g(t)\coloneqq 1+\lVert\sigma_{N}\rVert_{H^{1}}^{2},\quad h(t)\coloneqq 1.$
This then leads to the following double-exponential bound
$\sup_{t\in[0,T]}\biggl[\|\nabla u_{N}\|_{L^{2}}^{2}+\mathcal{E}^{-1}\|\nabla\sigma_{N}\|_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\|\Delta\sigma_{N}\|_{L^{2}}^{2}\biggr]\leq C\exp(\exp(CT)),$ (44)
for some $C$ independent of $N,\epsilon$, and $T$, but depending on $\mathcal{E},\alpha$, and the initial data. Note that the double-exponential nature of the bound comes from the application of the logarithmic Grönwall inequality to the single-exponential $L^{2}$-estimate (32).
$H^{2}$-estimate. One can compute directly that
$\nabla\mathcal{D}_{\epsilon}^{N}$ $=\frac{D(u_{N}):\big[\nabla D(u_{N})\big]}{\mathcal{D}_{\epsilon}^{N}}=\frac{\frac{1}{2}\nabla\lvert D(u_{N})\rvert^{2}}{\mathcal{D}_{\epsilon}^{N}},$ (45)